OpenStudy (anonymous):

Solve the following equations in the interval [0, 2π]. (Round your answers to 3 decimal places tan(x)sin(2x) = 1

OpenStudy (anonymous):

\[\tan(x)=\frac{sin(x)}{cos(x)}\] \[sin(2x)=2sin(x)cos(x)\] use these

OpenStudy (anonymous):

sin x * 2 sinx cos x = 1 ---- cos x 2sin^2 (x) cos x = cos x cos x (2sin^2 x - 1) = 0 cos x = 0 or 2sin^2 x = 1 can you proceed from here?

OpenStudy (anonymous):

i cant im sry i dont really get this

OpenStudy (anonymous):

if cos x = 0 you can find x using your calculator by entering 0 and pressing either cos-1 key or it might have arcsin key x will be pi/2 , 3pi/2 radians sin x^2 = 1/2 sin x = sqrt(1/2) or - sqrt (1/2) x will be pi/4, 3pi/4, 5pi/4 and 7pi/4 radians total solution is pi/2 , 3pi/2 , pi/4, 3pi/4, 5pi/4 and 7pi/4 radians

OpenStudy (anonymous):

thank you very much

OpenStudy (anonymous):

thats ok hope you've learned something